OpenStudy (anonymous):

3 men and 7 women can do a job in a 5 days, while4 men and 6 women can do it in 4 days. the number of days required for group of 10 women working together, as the same rate as before, to finish the same job is?

ganeshie8 (ganeshie8):

try to find a relationship between men and women work

ganeshie8 (ganeshie8):

3 men and 7 women can do a job in a 5 days, while4 men and 6 women can do it in 4 days total amount of work in man-women days : \((3m + 7w)*5\) which is also equal to : \((4m+6w)*4\)

ganeshie8 (ganeshie8):

equate them and find \(w\) in terms of \(m\)

ganeshie8 (ganeshie8):

amount of work is same, so \((3m + 7w)*5 = (4m+6w)*4\) \(\implies 15m + 35w = 16m + 24w\) \(\implies m = 11w\)

OpenStudy (anonymous):

10 women will do the same work in how many days?

ganeshie8 (ganeshie8):

you knw that 1 man = 11 women. use that relation and try to find how much time it takes for 10 women to finish the job

OpenStudy (anonymous):

okey

ganeshie8 (ganeshie8):

give it a try :)

ganeshie8 (ganeshie8):

since \(3m + 7w\) complete the job in \(5\) days, \(\implies 3*11w + 7w\) complete the job in \(5\) days you finish the rest ok ?

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np, let me knw wat do u get for answer :)

OpenStudy (anonymous):

10 women will do the same work in 20 days

ganeshie8 (ganeshie8):

since \(3m + 7w\) complete the job in \(5\) days, \(\implies 3*11w + 7w\) complete the job in \(5\) days \(\implies 40\) women complete the job in \(5\) days \(\implies 10\) women complete the job in \(20\) days

ganeshie8 (ganeshie8):

Perfect ! good job !!

OpenStudy (anonymous):

thanks thanku so mch

ganeshie8 (ganeshie8):

u wlc :)